OpenStudy (anonymous):

Solve the compound inequality 3x + 6 > 36 and 3x + 6 <= 54. Please show work? Thanks so much

OpenStudy (hoblos):

3x + 6 > 36 3x > 30 x>10 3x + 6 <= 54 3x <= 48 x<= 16 10 < x <=16

OpenStudy (dape):

Solve them separately, and then combine the solutions: \[ 3x+6-6 > 36-6 \\ 3x > 30 \\ x > 10 \] and \[ 3x+6-6 <= 54-6 \\ 3x <= 48 \\ x <= 16 \] So \[ 10 < x <= 16 \] Or on interval form: \[ x=(10,16] \]

OpenStudy (hoblos):

@Luis Rivera it is opened at 10 **